# latexText page

### v1

@@ 3.1 Preliminary on CLIP
The model is trained using contrastive learning on the matching scores, regarding the ground-truth image-text pairs as positive samples and other mismatched pairs as negatives. In practice, both encoders can be implemented as transformers. In §3.2 our method is derived based on the transformer architecture, and thus is suitable for interpreting both ViT-based image and transformer-based text encoders. Alternatively, the image encoder could be a CNN-based ResNet followed by an attention pooling layer, which is basically the same as an attention layer in the Transformer. Thus the proposed Grad-ECLIP is also applicable to the CNN-based CLIP, for which we display the visualization results for CLIP with ResNet [65] backbone in §4.1.5.
3.2 Methodology of Grad-ECLIP
@@
As shown in the illustration of Fig. 2, looking closely into the last layer of the network, the image embedding from visual encoder can be formulated as:
@@
where the approximation is based on assuming linearity of $\mathcal{LP}$. Noting that $\mathcal{LP}(x_{cls}^{(t)})=\mathcal{LP}(o_{cls}^{(t)}+x_{cls}^{(t+1)})$, we substitute recursively to obtain the approximation:
@@
and $\mathcal{A}$ represents the attention layer in the Transformer, $q_{cls}$ is the query embedding for the class token, while $k_{i}$ and $v_{i}\in\mathbb{R}^{C}$ represent the key and value embeddings at spatial location $i$, with $C$ as their channel dimension. [^1] We skip the superscript $(0)$ on $\{q,k,v\}$ identifying the layer for brevity. The $\mathrm{softmax}$ operation inside the attention layer measures the weight of the value on each location. Multi-heads are usually used in the attention layer to group the channel of $\{q,k,v\}$ into several heads, and (7) is operated inside each head with the $\mathrm{softmax}$ calculated over subsets of the channels. Then the final attention layer output is obtained by concatenating the results of each head together. In practice for visualization, we formulate the $o_{cls}^{(0)}$ with one attention head in the forward pass and operate the $\mathrm{softmax}$ over all channels as in (7). We discuss the influence multi-heads to visual explanation in the §4.3.3.
@@
where $v_{ic}$ is the c-th channel of $v_{i}$. Regarding $[v_{ic}]$ as the intermediate feature map, and $\lambda_{i}=\mathrm{softmax}(\tfrac{q_{cls}k_{i}^{\mathsf{T}}}{\sqrt{C}})$, then the importance of the $i$-th spatial location is defined as:
@@
The weight $w_{c}$ represents the importance of each feature channel, and $\lambda_{i}$ represents the importance of the values at each location via the softmax attention. However, from the visualization, we discover that the output of the $\mathrm{softmax}$ self attention function is extremely sparse. Important information may be encoded in different locations, but the $\mathrm{softmax}$ only selects the largest activation, which is not appropriate as a spatial weight. Therefore, we replace the $\mathrm{softmax}$ with a “loosened” correlation by applying 0-1 normalization on the similarities $[q_{cls}k_{i}^{\mathsf{T}}]_{i}$, i.e., $\lambda_{i}\approx\varPhi(q_{cls}k_{i}^{\mathsf{T}})$, where $\varPhi$ is the 0-1 normalization function applied over the set of similarities. In the experiments and §4.3.1, we compare using the loosened $\lambda_{i}$ and without $\lambda_{i}$ to show the effect of spatial weights, qualitatively and quantitatively.
@@
Visual and textual explanations: For the image encoder, the flattened heat map $H_{i}$ is reshaped and interpolated to the original image’s height and width, while for text encoder, the heat (importance) value on the $i_{th}$ tokens is remapped to the original word position in the sentence. Finally, based on the approximation in (9), the final explanation can be aggregated over all the layers by recursively processing each layer to obtain its heat map from (19). In the experiments, we use the last layer to explain the image encoder, and the last eight layers for interpreting the text encoder. The ablation study for the influence of different number of layers involved in image and text explanation is shown in §4.3.2.
CNN-based CLIP: Our proposed Grad-ECLIP can also be applied to CNN-based CLIP. The CNN-based CLIP model is composed of a ResNet backbone followed by an attention pooling. Thus, we can use the final attention layer in the pooling to conduct our explanation, which uses the same implementation as for ViT-based CLIP. The visualizations shown in §4.1.5 verify the effectiveness of Grad-ECLIP on CNN-based CLIP model.
@@
Figure 4: Explanations for image-text pairs from MS COCO using: by (a) raw self-attention; Transformer interpretation methods (b) Rollout, (c) GAME; and our method (d) Grad-ECLIP. The importances of words are visualized by the degree of green color.
4 Experiments with Grad-ECLIP
In this section we conduct experiments on Grad-ECLIP to: 1) evaluate its visual explanation qualitatively and quantitatively, and compare with the current SOTA methods; 2) evaluate the processing time; 3) conduct ablation studies, including the effect of spatial weight, the involved layers and attention heads.
Unless otherwise specified, we conducted the experiments with the ViT-B/16 architecture. We compared with representative baseline XAI methods from each category: 1) attention map-based Rollout [14], which takes into account all the attention maps computed along the forward pass, and raw attention in the last visual encoder layer, both of which are not result-specific explanation; 2) classical gradient-based method Grad-CAM [20], which takes the image-text similarity as target and calculate the gradients w.r.t. the ViT layer output; 3) relevance-based GAME [16], which integrates the relevancies and gradients propagated through the network; 4) cosine-based MaskCLIP [19] and CLIPSurgery [18], which generates similarity value on each location by the cosine between text feature and processed values as local image features; 5) M2IB [22], which applies information bottleneck principle to generate explanation map for CLIP. Each baseline is built with different properties and assumptions over the architecture. We also show visualization comparisons with the typical black-box perturbation method RISE [21], but did not conduct quantitative comparisons with black-box perturbation and Shapely methods, due to their computational complexity and inherent differences with our proposed approach.
@@ 4.1 Qualitative evaluation
In this section, we evaluate the proposed Grad-ECLIP qualitatively. The comparisons of visual explanations from different types of methods are presented in §4.1.1, and the visual explanations on both image and text encoder with image-sentence pair examples are preseneted in §4.1.2. Then, we conduct the visualizations of explanations on diverse image domains in §4.1.3 and adapt the Grad-ECLIP to other Vision Language Models (VLMs) in §4.1.4. Finally, we show that the Grad-ECLIP is applicable to CNN-based CLIP with ResNet backbone in §4.1.5.
4.1.1 Comparison of visual explanations
@@
We compare the visualizations of raw self-attention, Rollout, Grad-CAM, GAME, MaskCLIP, CLIPSurgery, M2IB, RISE and our Grad-ECLIP in Fig. 3 with images from ImageNet [66] and MS COCO [67]. Except raw attention and Rollout, which are defined to be text-agnostic, the others are all text-specific, so we test the same image with two different text inputs on MS COCO. Our Grad-ECLIP demonstrates a strong ability of generating clear and distinct text-specific heat maps, and gives reasonable explanation of verbs for interpreting CLIP. For example, the highlights for “holding” focus around the person’s hands (the 5th row of Fig. 3i), while “standing” highlights the person’s legs (the 6th row of Fig. 3i). We also notice that the sticks in the background are highlighted, which is probably because the sticks are regarded as “standing” in the snow.
@@
In contrast to our methods, Grad-CAM and MaskCLIP can produce highlights on the explained object, but both also generate significant noise. CLIPSurgery tends to put high and coarse attention on the object region, but also contains background noises. M2IB and MaskCLIP fail when the texts are verbs (“holding” and “standing”), while RISE performs the worst with interpreting CLIP model. The results of GAME and Rollout, which are both based on self-attentions of the model, generate confusing heat maps due to the sparse attention between tokens in some layers.
4.1.2 Explanations on image-sentence pairs
@@
The explanation map from Grad-ECLIP can also be generated from text encoder viewpoint. Using the gradient of matching score and the feature embeddings of word tokens, Grad-ECLIP can show the importance of each word in the given sentence when matching with an image. Fig. 4 shows example visual and textal explanations for image-text pairs from MS COCO. Although Rollout and GAME can highlight important words in the sentence, Grad-ECLIP is the only one showing good correspondence between image attention regions and important words. From the explanation of the sentence, we can identify which words are more important for CLIP when matching with the specific image, and correspndingly the text-specific important regions on the image are shown in the visual explanation. This word importance visualization of the input text can be helpful when designing text prompts for image-text dual-encoders in practical applications.
4.1.3 Visualization examples on diverse image domains
@@
We show the visualization comparison of different methods on the samples from different image domains, including the original ImageNet and ImageNet in different domains: rendition (ImageNet-R [68]), pencil sketch (ImageNet-Sketch [69]), natural adversarial example (ImageNet-A [70]), web images with captions (Conceptual Captions (CC) [71]), and chest x-ray with text (MSCXR [72]) in Fig. 5. For the image-caption pairs from the web-collected CC and chest X-ray data MSCXR, we generate explanations for both image and text encoder, and compare with the other methods that also provide text encoder explanations, including the raw attention, Rollout, GAME, M2IB.
@@
Our Grad-ECLIP explanations provide interesting insights into how CLIP handles different image domains. In Fig. 5 (top), given a normal banana image and text “banana”, Grad-ECLIP reveals that the yellow color is dominant to CLIP. However, when given a pencil sketch without color (ImageNet-Sketch), Grad-ECLIP reveals that CLIP looks at the curvature of the banana. For the color sketch of the banana (ImageNet-R), Grad-ECLIP shows that the color of the banana is mainly used, and not the black curved lines. Thus, from these examples, we may infer that CLIP prefers using the yellow color over the curved shape for matching with the “banana” text. With the sample of a web image and caption in the CC dataset, Grad-ECLIP generates clear and reasonable visual and the corresponding textual explanation map, showing that “cowboy”, “horses” and “desert” are the main used concepts in the matching (from high to low importance).
@@
Grad-ECLIP also provides interesting insights on how the original CLIP fails on novel domains. The last 2 rows of Fig. 5 show the explanations for chest x-ray images and text for the OpenAI CLIP model and a fine-tuned CLIP model (on MSCXR). The Grad-ECLIP explanation shows that the original CLIP uses the whole lobe to match with the words “defined” and “lobe”. In contrast, the fine-tuned CLIP locates the actual anomaly and matches it with the text “defined opacities largely”. The reason is that the fine-tuned model is trained to the specific domain that matches the X-ray and the illness location descriptions, while the original OpenAI CLIP model is more general and apparently “lobe” is the key word and the main object in the image-text pair.
@@
4.1.4 Explaining ViT classifier and BLIP with Grad-ECLIP
Since our explanation method is designed for CLIP encoders, which are Transformer-based, our method can be easily adapted to generate visual explanations for other Transformer-based models. Here we adapt Grad-ECLIP to ViT-based classifier [73] and BLIP [8] to show the generality of our method.
For explaining ViT-based classifier, the classification score on the corresponding category is used to calculate the gradient and generate the heat map. From the examples for ViT classifier in Fig. 6, we can see that although there are some noises on the background, Grad-ECLIP can well mark out the important region on the image for the specific class. As for the VLMs, shown by the visual explanation results presented in Fig. 6, when matching the same image-text pair, different models put attention on different regions. For example, BLIP notes the fins, while CLIP notes the fish body to match the image to “tench”. When matching with the sentence “a dog is playing with a frisbee”, BLIP puts more importance on the dog on the image, while CLIP shows places more importance on the frisbee.
Other VLMs like ALBEF [74] add additional attention layers after the encoders to fuse the image and text features, and thus our current method is not directly applicable since our method assumes that the last layer attention output has linear relationship with the final feature embedding. Our future work will investigate adapting our method to these modified ViT frameworks, e.g., the ALBEF model that uses cross attention to fuse image and text. Nonetheless, our ability to explain CLIP and other VLMs with similar architecture is significant considering that CLIP is by far the most widely used VLM.
@@ 4.1.4 Explaining ViT classifier and BLIP with Grad-ECLIP
4.1.5 Applying Grad-ECLIP to explain CNN-based CLIP

### v2

@@ III-A Preliminaries
The model is trained using contrastive learning on the matching scores, regarding the ground-truth image-text pairs as positive samples and other mismatched pairs as negatives. In practice, both encoders can be implemented as transformers. In §III-B, our method is derived based on the transformer architecture, and thus is suitable for interpreting both ViT-based image and transformer-based text encoders. Alternatively, the image encoder could be a CNN-based ResNet followed by an attention pooling layer, which is basically the same as an attention layer in the Transformer. Thus, the proposed Grad-ECLIP is also applicable to the CNN-based CLIP, for which we display the visualization results for CLIP with ResNet [18] backbone in §IV-A5.
III-B Methodology of Grad-ECLIP
@@
As shown in the illustration of Fig. 2, looking closely into the last layer of the network, the image embedding from the visual encoder can be formulated as:
@@
where the approximation is based on assuming linearity of $\mathcal{LP}$[45, 51]. Noting that $\mathcal{LP}(x_{cls}^{(t)})=\mathcal{LP}(o_{cls}^{(t)}+x_{cls}^{(t+1)})$, we substitute recursively to obtain the approximation:
@@
and $\mathcal{A}$ represents the attention layer in the Transformer, $q_{cls}$ is the query embedding for the class token, while $k_{i}$ and $v_{i}\in\mathbb{R}^{C}$ represent the key and value embeddings at spatial location $i$, with $C$ as their channel dimension. [^1] We skip the superscript $(0)$ on $\{q,k,v\}$ identifying the layer for brevity. The $\mathrm{softmax}$ operation inside the attention layer measures the weight of the value on each location. Multi-heads are usually used in the attention layer to group the channels of $\{q,k,v\}$ into several heads, and (7) is operated inside each head with the $\mathrm{softmax}$ calculated over subsets of the channels. Then, the final attention layer output is obtained by concatenating the results of each head together. In practice for visualization, we formulate the $o_{cls}^{(0)}$ with one attention head in the forward pass and operate the $\mathrm{softmax}$ over all channels as in (7). We discuss the influence of multi-heads on visual explanation in the §IV-C3.
@@
where $v_{ic}$ is the c-th channel of $v_{i}$. Regarding $v_{ic}]$ as the intermediate feature map, and $\lambda_{i}=\mathrm{softmax}(\tfrac{q_{cls}k_{i}^{\mathsf{T}}}{\sqrt{C}})$, then the importance of the $i$-th spatial location is defined as:
@@
The weight $w_{c}$ represents the importance of each feature channel, and $\lambda_{i}$ represents the importance of the values at each location via the softmax attention. However, from the visualization, we discover that the output of the $\mathrm{softmax}$ self-attention function is extremely sparse. Important information may be encoded in different locations, but the $\mathrm{softmax}$ only selects the largest activation, which is not appropriate as a spatial weight. Therefore, we replace the $\mathrm{softmax}$ with a “loosened” correlation by applying 0-1 normalization on the similarities $q_{cls}k_{i}^{\mathsf{T}}]_{i}$, i.e., $\lambda_{i}\approx\varPhi(q_{cls}k_{i}^{\mathsf{T}})$, where $\varPhi$ is the 0-1 normalization function applied over the set of similarities. In the experiments and §IV-C1, we compare using the loosened $\lambda_{i}$ and without $\lambda_{i}$ to show the effect of spatial weights, qualitatively and quantitatively.
@@
Visual and textual explanations: For the image encoder, the flattened heat map $H_{i}$ is reshaped and interpolated to the original image’s height and width, while for the text encoder, the heat (importance) value on the $i_{th}$ tokens is remapped to the original word position in the sentence. Finally, based on the approximation in (9), the final explanation can be aggregated over all the layers by recursively processing each layer to obtain its heat map from (19). In the experiments, we use the last layer to explain the image encoder, and the last eight layers for interpreting the text encoder. The ablation study for the influence of different numbers of layers involved in image and text explanation is shown in §IV-C2.
CNN-based CLIP: Our proposed Grad-ECLIP can also be applied to CNN-based CLIP. The CNN-based CLIP model is composed of a ResNet backbone followed by an attention pooling. Thus, we can use the final attention layer in the pooling to conduct our explanation, which uses the same implementation as for ViT-based CLIP. The visualizations shown in §IV-A5 verify the effectiveness of Grad-ECLIP on CNN-based CLIP model.
@@
Figure 4: Explanations for image-text pairs from MS COCO using: by (a) raw self-attention; Transformer interpretation methods (b) Rollout, (c) GAME; and our method (d) Grad-ECLIP. The importance of words is visualized by the degree of green color.
IV Experiments with Grad-ECLIP
In this section, we conduct experiments on Grad-ECLIP to: 1) evaluate its visual explanation qualitatively and quantitatively, and compare with the current SOTA methods; 2) evaluate the processing time; 3) conduct ablation studies, including the effect of spatial weight, the involved layers, and attention heads.
Unless otherwise specified, we conducted the experiments with the ViT-B/16 architecture. We compared with representative baseline XAI methods from each category: 1) attention map-based Rollout [1], which takes into account all the attention maps computed along the forward pass, and raw attention in the last visual encoder layer, both of which are not result-specific explanations; 2) classical gradient-based method Grad-CAM [46], which takes the image-text similarity as target and calculates the gradients w.r.t. the ViT layer output; 3) relevance-based GAME [9], which integrates the relevancies and gradients propagated through the network; 4) cosine-based MaskCLIP [73] and CLIPSurgery [29], which generate similarity value on each location by the cosine between text feature and processed values as local image features; 5) M2IB [58], which applies the information bottleneck principle to generate an explanation map for CLIP. Each baseline is built with different properties and assumptions over the architecture. For evaluation of text encoder explanations, we compare Rollout, GAME, and M2IB, and our Grad-ECLIP by generating word importance on the input sentences – the other methods (Grad-CAM, MaskCLIP, and CLIPSurgery) are excluded since they are only applicable to the image encoder. We also show visualization comparisons with the typical perturbation method RISE [36] (and IG [50] in the Appendix), but did not conduct quantitative comparisons with perturbation and Shapley methods, due to their computational complexity and black-box property. Since the proposed Grad-ECLIP is a white-box method designed for CLIP, we prioritize comparison with other XAI methods designed for CLIP (CLIPSurgery, MaskCLIP, M2IB) or similar transformer architectures (Rollout, GAME), and the classical CAM-based method Grad-CAM, both qualitatively and quantitatively.
@@ IV-A Qualitative evaluation
In this section, we evaluate the proposed Grad-ECLIP qualitatively. The comparisons of visual explanations from different types of methods are presented in §IV-A1, and the visual explanations on both image and text encoders with image-sentence pair examples are presented in §IV-A2. Then, we conduct the visualizations of explanations on diverse image domains in §IV-A3 and adapt the Grad-ECLIP to other Vision Language Models (VLMs) in §IV-A4. Finally, we show that the Grad-ECLIP is applicable to CNN-based CLIP with ResNet backbone in §IV-A5.
IV-A1 Comparison of visual explanations
@@
We compare the visualizations of raw self-attention, Rollout, Grad-CAM, GAME, MaskCLIP, CLIPSurgery, M2IB, RISE, and our Grad-ECLIP in Fig. 3 with images from ImageNet [43] and MS COCO [31]. Except for raw attention and Rollout, which are defined to be text-agnostic, the others are all text-specific, so we test the same image with two different text inputs on MS COCO. Our Grad-ECLIP demonstrates a strong ability of generating distinct text-specific heat maps, and gives corresponding explanation of verbs for interpreting CLIP. For example, the highlights for “holding” focus around the person’s hands (the 5th row of Fig. 3i), while “standing” highlights the person’s legs (the 6th row of Fig. 3i). We also notice that the sticks in the background are highlighted, which is probably because the sticks are regarded as “standing” in the snow.
@@
In contrast to our methods, Grad-CAM and MaskCLIP can produce highlights on the explained object, but both also generate significant noise. CLIPSurgery tends to put high and coarse attention on the object region, but also contains background noise. M2IB and MaskCLIP fail when the texts are verbs (“holding” and “standing”), while RISE performs the worst when interpreting the CLIP model. The results of GAME and Rollout, which are both based on the self-attention of the model, generate confusing heat maps due to the sparse attention between tokens in some layers.
IV-A2 Explanations on image-sentence pairs
@@
The explanation map from Grad-ECLIP can also be generated from the text encoder viewpoint. Using the gradient of the matching score and the feature embeddings of word tokens, Grad-ECLIP can show the importance of each word in the given sentence when matching with an image. Fig. 4 shows example visual and textual explanations for image-text pairs from MS COCO. Although Rollout and GAME can highlight important words in the sentence, Grad-ECLIP is the only one showing good correspondence between image attention regions and important words. From the explanation of the sentence, we can identify which words are more important for CLIP when matching with the specific image, and correspndingly the text-specific important regions on the image are shown in the visual explanation. This word importance visualization of the input text can be helpful when designing text prompts for image-text dual-encoders in practical applications.
IV-A3 Visualization examples on diverse image domains
@@
We show the visualization comparison of different methods on the samples from different image domains, including the original ImageNet, ImageNet in different domains: rendition (pencil sketch (ImageNet-Sketch [55]), natural adversarial example (ImageNet-A [19]), web images with captions (Conceptual Captions (CC) [47]), and chest x-ray with text (MSCXR [4]) in Fig. 5. For the image-caption pairs from the web-collected CC and chest X-ray data MSCXR, we generate explanations for both the image and text encoders, and compare with the other methods that also provide text encoder explanations, including the raw attention, Rollout, GAME, and M2IB.
@@
Our Grad-ECLIP explanations provide interesting insights into how CLIP handles different image domains. In Fig. 5 (top), given a normal banana image and text “banana”, Grad-ECLIP reveals that the yellow color is dominant to CLIP. However, when given a pencil sketch without color (ImageNet-Sketch), Grad-ECLIP reveals that CLIP looks at the curvature of the banana. For the color sketch of the banana (ImageNet-R), Grad-ECLIP shows that the color of the banana is mainly used, and not the black curved lines. Thus, from these examples, we may infer that CLIP prefers using the yellow color over the curved shape for matching with the “banana” text. With the sample of a web image and caption in the CC dataset, Grad-ECLIP generates visual highlights and the corresponding textual explanation map, showing that “cowboy”, “horses” and “desert” are the main concepts used in the matching (from high to low importance).
@@
Grad-ECLIP also provides interesting insights into how the original CLIP fails on novel domains. The last 2 rows of Fig. 5 show the explanations for chest x-ray images and text for the OpenAI CLIP model and a fine-tuned CLIP model (on MSCXR). The Grad-ECLIP explanation shows that the original CLIP uses the whole lobe to match with the words “defined” and “lobe”. In contrast, the fine-tuned CLIP locates the actual anomaly and matches it with the text “defined opacities largely”. The reason is that the fine-tuned model is trained to the specific domain that matches the X-ray and the illness location descriptions, while the original OpenAI CLIP model is more general, and apparently “lobe” is the keyword and the main object in the image-text pair.
@@
IV-A4 Explaining ViT classifier and BLIP with Grad-ECLIP
Since our explanation method is designed for CLIP encoders, which are Transformer-based, our method can be easily adapted to generate visual explanations for other Transformer-based models. Here we adapt Grad-ECLIP to a ViT-based classifier [12] and BLIP [26] to show the generality of our method.
For explaining the ViT-based classifier, the classification score on the corresponding category is used to calculate the gradient and generate the heat map. From the examples for the ViT classifier in Fig. 6, we can see that although there is some noise on the background, Grad-ECLIP can well mark out the important region on the image for the specific class. As for the VLMs, shown by the visual explanation results presented in Fig. 6, when matching the same image-text pair, different models put attention on different regions. For example, BLIP notes the fins, while CLIP notes the fish body to match the image to “tench”. When matching with the sentence “a dog is playing with a frisbee”, BLIP puts more importance on the dog in the image, while CLIP places more importance on the frisbee.
Other VLMs like ALBEF [27] add additional attention layers after the encoders to fuse the image and text features, and thus our current method is not directly applicable since our method assumes that the last layer attention output has a linear relationship with the final feature embedding. Our future work will investigate adapting our method to these modified ViT frameworks, e.g., the ALBEF model that uses cross attention to fuse image and text. Nonetheless, our ability to explain CLIP and other VLMs with similar architecture is significant considering that CLIP is by far the most widely used VLM.
@@ IV-A4 Explaining ViT classifier and BLIP with Grad-ECLIP
IV-A5 Applying Grad-ECLIP to explain CNN-based CLIP